40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jan 25 20:42:46 2011, Osamu, Update, Green Locking, Slow servo for green laser 
    Reply  Tue Jan 25 21:57:59 2011, Koji, Update, Green Locking, Slow servo for green laser 110125_Xend_thermal.pdf
Message ID: 4201     Entry time: Tue Jan 25 20:42:46 2011     Reply to this: 4202
Author: Osamu 
Type: Update 
Category: Green Locking 
Subject: Slow servo for green laser 

I implemented a slow servo for green laser thermal control on c1scx.mdl. Ch6,7 of ADC and ch6 of DAC are assigned for this servo as below;

 

Ch6 of ADC: PDH error signal

CH7 of ADC: PZT feedback signal

CH6 of DAC: feedback signal to thermal of green laser

 

Note that old EPICS themal control cable is not hooked anymore.

I made a simple MEDM screen(...medm/c1scx/master/C1SCX_BCX_SLOW.adl) linked from GREEN medm screen (C1GCV.adl) on sitemap.

During this work, I noticed that some of the epics switch is not recovered by autoburt. What I noticed is filter switch of SUSPOS, SUSPIT, SUSYAW, SDSEN, and all coil output for ETMX.

I had no idea to fix them, probably Joe knows. I guess other suspensitons has the same problems.

ELOG V3.1.3-